40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 13 03:18:14 2011, kiwamu, Update, LSC, Y arm locking prep image.png
    Reply  Tue Sep 13 19:31:09 2011, kiwamu, Update, LSC, Y arm locking prep Yarmlocked.png
       Reply  Wed Sep 14 01:17:51 2011, kiwamu, Update, LSC, measurement of spot position on Y arm misposition.png
          Reply  Wed Sep 14 15:00:00 2011, kiwamu, Update, LSC, measurement of spot position on Y arm DSC_3408_small.jpg
    Reply  Wed Sep 14 21:18:45 2011, kiwamu, Update, LSC, Y arm locking prep 
Message ID: 5410     Entry time: Wed Sep 14 21:18:45 2011     In reply to: 5392
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: Y arm locking prep 

Although we did some of the Input Matrix diagonalization, we have not yet actually used this knowledge. As a result all of the optics are shaking all over the place.

Sunshine Task: Set the input matrices to their calculated values and then adjust the OSEM damping gains for all optics so as to get a Q ~ 5.

ELOG V3.1.3-